Find the derivatives of this function


In(ye^x)=3

Grade:

2 Answers

Pawan Kumar Sharma
12 years ago

Dear ilham,

ln a = b, then e^b = a

similarly, ln(ye^x) = 3 => e^3 = ye^x

y =  e^(3-x)

differentiating on both sides,

dy/dx = -e^(3-x)

Plz approve the answer.....

Veenus Yadav
12 years ago

     in(yex)=3                                                                              y=e3/ex

=>yex=e3

=>(yex )'=(e3)'

=>yex+exdy/dx=0

=>dy/dx=-y =-e3-x